OpenStudy (anonymous):

prove identity: (1+sinx)/(1-sinx)=([1/sinx]+1)/([1/sinx]-1)

OpenStudy (anonymous):

starting with right hand side 1 -- + 1 = 1 + sin x sin x -------- sin x 1 -- - 1 = 1 - sinx sin x ------- sin x dividng first exoression by the last, we invert the last one and multiply: so this is 1 + sin x * sin x -------- ------ sin x 1 - sin x = (1 + sin x ) / ( 1 - sin x)

OpenStudy (anonymous):

...Ohh i got lost. Can you do it step by step please?

OpenStudy (anonymous):

the rhs is 1/sinx+1/1/sinx-1, so it becomes 1+sinx/sinx?

OpenStudy (anonymous):

no (1/sinx )+ 1 = (1 + sinx) / sinx and ( 1 /sinx) - 1 = (1 - sinx) / sin x

OpenStudy (anonymous):

and dividing the 2 above results gives the left side

OpenStudy (anonymous):

OHHHH OMGOSH I GET IT!

OpenStudy (anonymous):

if the two fractions were 3 / 4 and 4 / 5 then i to work out 3/4 / 4/5 we change 4/5 to 5/4 and then multiply 3/4 * 5/4 = 15/16 if the numbers are replaced by terms in x or sinx etc the same rules apply

OpenStudy (anonymous):

ohs! i see

OpenStudy (anonymous):

good!!

OpenStudy (anonymous):

can you help me with another question..i kind of got stuck = u =

OpenStudy (anonymous):

[(1/cosx)+tanx](1-sinx)=cosx i know i have to work on the LHS and i don't know if i should put tanx to sinx/cosx or change 1/cosx to secx